# Struct2D: A Perception-Guided Framework for Spatial Reasoning in Large Multimodal Models

### v1

@@ Appendix B Details of Struct2D-Set
Overview. Struct2D-Set is a large-scale instruction tuning dataset aimed at enabling spatial reasoning and scene understanding in indoor 3D environments using only 2D projected inputs. It contains over 200K question-answer (QA) pairs derived from 6K richly annotated indoor scenes drawn from ScanNet [19], ScanNet++[87], and ARKitScenes[4]. Each QA instance is paired with structured scene- and object-level metadata, allowing models to learn spatial concepts without relying on explicit 3D feature representations during training.
@@
Spatial Relationship. This category evaluates a model’s ability to reason about the directional relationships between objects in a 3D scene from an egocentric perspective. Following the formulation in VSI-Bench [81], we focus on the subtask of relative direction, where the goal is to identify the directional location of a target object based on a specified standing point and facing direction.
@@
Other Categories. The remaining task types—such as object attribute identification and binary attribute verification—are constructed by augmenting QA pairs from existing 3D vision-language datasets, including ScanQA [3], SQA3D [52], Scan2Cap [14], ScanRefer [10], and Multi3DRefer [94]. These datasets provide scene-specific questions grounded in the ScanNet environment and collectively cover all eight spatial reasoning categories defined in Struct2D-Set.
To adapt these examples for instruction tuning, we first use ChatGPT to rephrase each question into a more natural and instructional style. For tasks requiring reasoning, we also prompt ChatGPT to generate long-form answers that walk through the inference process. For visual grounding, we localize referenced objects using two approaches: when object IDs are available, we apply our keyframe detection method to extract representative egocentric views. In datasets with descriptive references (e.g., SQA3D), we extract referring expressions with ChatGPT and apply Grounding DINO [50] and SAM [35] to identify and segment the mentioned objects.
@@ Appendix C Implementation Details
We use Qwen2.5VL [71] as the base large multimodal model (LMM) for instruction tuning. During training, the model receives BEV images with filtered object marks and object-centric metadata as core inputs. For tasks requiring visual cues such as object color or quantity, we additionally provide egocentric keyframes. The BEV images are resized to $640\times 640$, with object marks adaptively scaled based on their original image resolution. Keyframes are resized to $256\times 246$ and stitched into compact 1×2 or 2×4 grids, enabling efficient batch loading and reducing GPU memory consumption.
@@
At evaluation time, we follow standard practices from prior work [30, 62], reconstructing point clouds offline using BundleFusion [18], detecting 3D objects using Mask3D[65] and UniDet [36], and projecting the results to produce BEV images and 2D object marks. For object-level grounding, we apply a rule-based method to identify the relevant objects mentioned in each question.
@@ Appendix D Results on 3D Grounding and 3D Dense Captioning
Table 6: 3D Grounding Evaluation on ScanRefer [10] and Multi3DRefer [94] datasets.
@@
3DVG-Transformer [96] 47.6 34.7 25.5
3DJCG [6] 49.6 37.3 26.6
D3Net [11] 37.9 32.2
M3DRef-CLIP [94] 51.9 44.7 42.8 38.4
@@
Chat-Scene [30] 55.5 50.2 57.1 52.4
@@
Qwen2-VL-7B [71] 5.4 5.1 21.1 19.9
Qwen2-VL-7B (GPT4Scene [62]) 40.5 36.7 45.4 42.1
@@
Table 7: 3D Dense Captioning Evaluation on Scan2Cap [14] dataset.
@@
Scan2Cap [14] 34.2 55.3 23.3 44.5
3DJCG [6] 40.2 59.2 31.0 50.8
X-Trans2Cap [89] 35.7 54.7 25.1 45.3
3D-VisTA [101] 36.5 57.6 34.0 54.3
Vote2Cap-DETR [12] 39.3 59.3 34.5 54.4
@@
LL3DA [13] 41.4 59.5 36.8 55.1
LEO [31] 36.9 57.8
Chat-Scene [30] 38.2 60.6 36.3 58.1
Robin3D [34] 38.4
@@
Qwen2-VL-7B [71] 3.8 24.7 3.8 24.6
Qwen2-VL-7B (GPT4Scene [62]) 36.3 57.6 34.2 55.2
@@
Quantitative results. Tables 6 and 7 present our model’s performance on 3D grounding (ScanRefer, Multi3DRefer) and dense captioning (Scan2Cap) benchmarks. While our method does not achieve the highest scores under rule-based metrics such as all F1@0.25/0.5 and BLEU/ROUGE, it consistently delivers competitive results compared to existing vision-language baselines. Importantly, our approach does not rely on point cloud features during training or evaluation, in contrast to task-specific and 3D LLM models that depend heavily on explicit 3D representations. In addition, our approach requires substantially fewer egocentric keyframes on average (2 compared to 8 in GPT4Scene [62]), resulting in a more efficient and scalable training process. Compared to models designed for narrow tasks, our framework is more general and supports a wider range of spatial reasoning types, including relative direction and route planning, which are not covered by these benchmarks. It is also worth noting that the current evaluation metrics are rule-based and limited in expressiveness, which may not fully reflect a model’s capability in spatial understanding.
@@
Appendix E Broader impacts
Our work introduces Struct2D, a perception-guided prompting framework that enables Large Multimodal Models (LMMs) to perform robust spatial reasoning in 3D environments using only structured 2D and text inputs. This direction offers several broader implications for research, society, and the reasonable development of AI systems.
@@ Appendix E Broader impacts
Research Contributions. By decoupling LMM training from explicit 3D input requirements, Struct2D promotes research into modular, scalable instruction-tuning pipelines that can generalize across environments with different sensor setups. Furthermore, our public release of Struct2D-Set—a large-scale spatial instruction dataset built with a principled blend of structured prompts, egocentric frames, and metadata—contributes valuable benchmarks to the broader vision-language community.

### v3

@@ Appendix B Details of Struct2D-Set
Overview. Struct2D-Set is a large-scale instruction tuning dataset aimed at enabling spatial reasoning and scene understanding in indoor 3D environments using only 2D projected inputs. It contains over 200K question-answer (QA) pairs derived from 6K richly annotated indoor scenes drawn from ScanNet [dai2017scannet], ScanNet++[yeshwanth2023scannet++], and ARKitScenes[baruch2021arkitscenes]. Each QA instance is paired with structured scene- and object-level metadata, allowing models to learn spatial concepts without relying on explicit 3D feature representations during training.
@@
Spatial Relationship. This category evaluates a model’s ability to reason about the directional relationships between objects in a 3D scene from an egocentric perspective. Following the formulation in VSI-Bench [yang2024thinking], we focus on the subtask of relative direction, where the goal is to identify the directional location of a target object based on a specified standing point and facing direction.
@@
Other Categories. The remaining task types—such as object attribute identification and binary attribute verification—are constructed by augmenting QA pairs from existing 3D vision-language datasets, including ScanQA [azuma2022scanqa], SQA3D [ma2022sqa3d], Scan2Cap [chen2021scan2cap], ScanRefer [chen2020scanrefer], and Multi3DRefer [Multi3DRefer]. These datasets provide scene-specific questions grounded in the ScanNet environment and collectively cover all eight spatial reasoning categories defined in Struct2D-Set.
To adapt these examples for instruction tuning, we first use ChatGPT to rephrase each question into a more natural and instructional style. For tasks requiring reasoning, we also prompt ChatGPT to generate long-form answers that walk through the inference process. For visual grounding, we localize referenced objects using two approaches: when object IDs are available, we apply our keyframe detection method to extract representative egocentric views. In datasets with descriptive references (e.g., SQA3D), we extract referring expressions with ChatGPT and apply Grounding DINO [liu2024grounding] and SAM [kirillov2023segment] to identify and segment the mentioned objects.
@@ Appendix C Implementation Details
We use Qwen2.5VL [wang2024qwen2] as the base multimodal large language model (MLLM) for instruction tuning. During training, the model receives BEV images with filtered object marks and object-centric metadata as core inputs. For tasks requiring visual cues such as object color or quantity, we additionally provide egocentric keyframes. The BEV images are resized to $640\times 640$, with object marks adaptively scaled based on their original image resolution. Keyframes are resized to $256\times 246$ and stitched into compact 1×2 or 2×4 grids, enabling efficient batch loading and reducing GPU memory consumption.
@@
At evaluation time, we follow standard practices from prior work [huang2023chat, qi2025gpt4scene], reconstructing point clouds offline using BundleFusion [dai2017bundlefusion], detecting 3D objects using Mask3D[schult2023mask3d] and UniDet [kolodiazhnyi2024unidet3d], and projecting the results to produce BEV images and 2D object marks. For object-level grounding, we apply a rule-based method to identify the relevant objects mentioned in each question.
@@ Appendix D Results on 3D Grounding and 3D Dense Captioning
Table 6: 3D Grounding Evaluation on ScanRefer [chen2020scanrefer] and Multi3DRefer [Multi3DRefer] datasets.
@@
3DVG-Transformer [zhao20213dvg] 47.6 34.7 25.5
3DJCG [3djcg] 49.6 37.3 26.6
D3Net [chen2021d3net] 37.9 32.2
M3DRef-CLIP [Multi3DRefer] 51.9 44.7 42.8 38.4
@@
Chat-Scene [huang2023chat] 55.5 50.2 57.1 52.4
@@
Qwen2-VL-7B [wang2024qwen2] 5.4 5.1 21.1 19.9
Qwen2-VL-7B (GPT4Scene [qi2025gpt4scene]) 40.5 36.7 45.4 42.1
@@
Table 7: 3D Dense Captioning Evaluation on Scan2Cap [chen2021scan2cap] dataset.
@@
Scan2Cap [chen2021scan2cap] 34.2 55.3 23.3 44.5
3DJCG [3djcg] 40.2 59.2 31.0 50.8
X-Trans2Cap [X-trans2cap] 35.7 54.7 25.1 45.3
3D-VisTA [zhu20233d] 36.5 57.6 34.0 54.3
Vote2Cap-DETR [vote2cap-detr] 39.3 59.3 34.5 54.4
@@
LL3DA [chen2024ll3da] 41.4 59.5 36.8 55.1
LEO [huang2023embodied] 36.9 57.8
Chat-Scene [huang2023chat] 38.2 60.6 36.3 58.1
Robin3D [Robin3D] 38.4
@@
Qwen2-VL-7B [wang2024qwen2] 3.8 24.7 3.8 24.6
Qwen2-VL-7B (GPT4Scene [qi2025gpt4scene]) 36.3 57.6 34.2 55.2
@@
Quantitative results. Tables 6 and 7 present our model’s performance on 3D grounding (ScanRefer, Multi3DRefer) and dense captioning (Scan2Cap) benchmarks. While our method does not achieve the highest scores under rule-based metrics such as all F1@0.25/0.5 and BLEU/ROUGE, it consistently delivers competitive results compared to existing vision-language baselines. Importantly, our approach does not rely on point cloud features during training or evaluation, in contrast to task-specific and 3D LLM models that depend heavily on explicit 3D representations. In addition, our approach requires substantially fewer egocentric keyframes on average (2 compared to 8 in GPT4Scene [qi2025gpt4scene]), resulting in a more efficient and scalable training process. Compared to models designed for narrow tasks, our framework is more general and supports a wider range of spatial reasoning types, including relative direction and route planning, which are not covered by these benchmarks. It is also worth noting that the current evaluation metrics are rule-based and limited in expressiveness, which may not fully reflect a model’s capability in spatial understanding.
@@
Appendix E Failure cases
Figure 13: Examples of failure cases caused by 3D reconstruction and detection.
To better understand the limitations of our approach, we conducted a qualitative error analysis on 30 representative questions spanning multiple QA types in VSI-Bench. Among the 16 failure cases, we identified two dominant causes. First, in 11 cases, the underlying 3D reconstruction was noisy or incomplete, producing degraded BEV projections that obscured critical spatial layouts. Second, in 5 cases, missing detections—often involving small or heavily occluded objects—led to incomplete structured inputs. Both factors reduce Struct2D’s ability to encode accurate spatial cues, thereby hindering its reasoning capability. Representative visualizations of these failure modes are shown in Figure 13. These examples highlight the importance of robust 3D reconstruction and reliable object grounding for spatial reasoning in complex indoor scenes.
Appendix F Broader impacts
Our work introduces Struct2D, a perception-guided prompting framework that enables Multimodal Large Language Models (MLLMs) to perform robust spatial reasoning in 3D environments using only structured 2D and text inputs. This direction offers several broader implications for research, society, and the reasonable development of AI systems.
@@ Appendix F Broader impacts
Research Contributions. By decoupling MLLM training from explicit 3D input requirements, Struct2D promotes research into modular, scalable instruction-tuning pipelines that can generalize across environments with different sensor setups. Furthermore, our public release of Struct2D-Set—a large-scale spatial instruction dataset built with a principled blend of structured prompts, egocentric frames, and metadata—contributes valuable benchmarks to the broader vision-language community.